For 2y+8x=-6 determine the value of y for both x = -3 and x = -1, then graph the line using these two points.
Answer by Alan3354(69443)

For 2y+8x=-6 determine the value of y for both x = -3 and x = -1, then graph the line using these two points.
---------
2y+8x=-6
y + 4x = -3
y = -4x - 3
============
Sub the values for x and find y.
========
Plot the 2 points, draw a line thru them.